Please help me solve this Geometric Sequence.
My teacher doesn't teach very well so it's kind of hard to comprehend everything.
equation-91st term: 1/5,-1/5,1/5,-1/5,1/5,...

Answer by drk(1908)

I hope I can help you here.
The first term, a1 = 1/5
the common ratio = -1.
----
From here not sure what you want:
91st term =1/5
formula is
Tn = (1/5)*(-1)^(n-1)
The infinite sum is a/(1-r) = (1/5) / (1-1/5) = 1/4
----
hopefully one of these is what you wanted